litboi:

2. Find the geometric mean of the pair of numbers. 36 and 4 (1 point)22 17 12 32 3. What are the values of a and b? (1 point)a = 8, b = 2√17 a = 18, b = 2√17 a = 8, b = 8√17 a = 64, b = 68 4. What is the value of x, given that ? (1 point)x = 10 x = 20 x = 13 x = 25.5 5. What is the value of x? (1 point)8 12 6 2 6. bisects NLM, LM = 18, NO = 4, and LN = 10. What is the value of x? (1 point)7.2 45 5.43 2.22 7. What is the value of x

563blackghost:

2. The geometric mean is found by multiply the pair then squaring the product. The rest of the questions are incomplete. Please do provide the needed information to solve them.

ratnakermehta20:

By definition, the Geometric Mean (GM) G of two positive numbers (mos.) x and y is \[G=\sqrt{xy}.\] Hence, the GM of 36 and 4 is \[\sqrt{36*4}=\sqrt{144}=12.\]